OpenStudy (anonymous):

Find the equation of a parallel to the line y=17x and contains the point (6,-5)

OpenStudy (anonymous):

parallel lines have the same slope sooooo we know the slope=17 use the equation y-y1=m(x-x1) y+5= 17(x-6) y+5=17x- 102 y=17x-107 is the answer

OpenStudy (anonymous):

Thank you so much I have been staring at this problem for days and nothing clicked thank you so much

OpenStudy (anonymous):

np:)